# Reinforcement Twinning: from digital twins to model-based reinforcement learning

### 31 Jan 2024 v2

@@ Abstract
We propose a novel framework for simultaneously training the digital twin of an engineering system and an associated control agent. The training of the twin combines methods from data assimilation and system identification, while the training of the control agent combines model-based optimal control and model-free reinforcement learning. The combined training of the control agent is achieved by letting it evolve independently along two paths: one driven by a model-based optimal control and another driven by reinforcement learning. The virtual environment offered by the digital twin is used as a playground for confrontation and indirect interaction. This interaction occurs as an “expert demonstrator”, where the best policy is selected for the interaction with the real environment and “cloned” to the other if the independent training stagnates. We refer to this framework as Reinforcement Twinning (RT).
@@ 1 Introduction
Nonlinear system identification is mostly concerned with the model identification problem. The system is subject to external inputs and observed through noisy measurements but is deterministic. The model can be fit into structures that are derived either from first principles or from general templates such as NRAX or Volterra models (Schoukens and Ljung, 2019), and the main goal is to predict how the system responds to actuation. The traditional context is control engineering, i.e. driving the system towards a desired trajectory. Besides the stronger focus on input-output relation (and thus the focus on actuated systems), the main difference with respect to traditional data assimilation is in the deterministic nature of the system to be inferred: errors in the initial conditions are not amplified as dramatically as, for example, in weather forecasting. Therefore, the resulting uncertainties remain within the bounds of model uncertainties. Moreover, in many actuated systems, the forced response dominates over the natural one and the initial conditions are quickly forgotten. Tools for system identification are reviewed by Nelles (2001) and Suykens et al. (1996).
@@
While machine learning enters the literature of data assimilation and system identification, cross-fertilization proceeds bilaterally. The need for identifying dynamic models from data is also growing in the literature on reinforcement learning, which is significantly drawing ideas from system identification and optimal control theory. Reinforcement learning is a subset of machine learning concerned with the training of an agent via trial and error to achieve a goal while acting on an environment (Sutton and Barto, 2018; Bertsekas, 2019). The framework differs from classic control theory because of its roots in sequential decision making: the environment to be controlled and the agent constitute a Markov Decision Process (MDP, Puterman (1994)) rather than a dynamical system. Model-free approaches solely rely on input-output information and function approximations of the agent’s performance measures, while model-based approaches combine these with a model of the environment (Moerland et al., 2022a; Luo et al., 2022a). The model can be used as a forecasting tool, thus playing the same role as in model predictive control, or as a playground to accelerate the learning (Schwenzer et al., 2021). Model-free approaches have gained burgeoning popularity thanks to their success in video games (Szita, 2012; Mnih et al., 2013a) or natural language processing (Uc-Cetina et al., 2022), for which the definition of a system model is cumbersome. However, these algorithms have shown significant and somewhat surprising limitations for engineering problems governed by simple PDEs (Werner and Peitz, 2023; Pino et al., 2023).
This article proposes a framework combining ideas from data assimilation, system identification and reinforcement learning with the goal of: (1) training a digital twin on real-time data and (2) solving a control problem. We refer to this framework as Reinforcement Twinning (RT). The rest of the article is structured as follows. Section 2 provides a concise literature review of related works across the intersected disciplines and identifies the main novelties of the proposed framework. Section 3 presents the general framework and the key definitions, while Section 4 goes into the mathematical details of its implementation. Section 5 introduces three test cases selected for their relevance in the engineering literature and stemming from largely different fields. These are(1) the control of a wind turbine subject to time-varying wind speed, (2) the trajectory control of flapping-wing micro air vehicles (FWMAVs) subject to wind gusts, and (3) the mitigation of thermal loads in the management of cryogenic storage tanks. Although all test cases are analyzed using synthetic data and are thus somewhat contrived, the main focus is here to show that the model-based and the model-free training can benefit from each other. Section 6 presents the main results; section 7 closes with conclusions and perspectives.
@@ 2.2 Machine learning for System Identification and Control
Important contributions within the first class of approaches are the works by Ljung et al. (2020) and Gonzalez and Yu (2018), who use LSTM networks (Hochreiter and Schmidhuber, 1997). These are variants of RNNs, which are very popular in natural language processing and use three gates (forget/input/output) to preserve information over longer time steps, thus allowing for better handling of long-term dependencies. Canaday et al. (2020) used ESNs, which gained popularity for forecasting in stochastic systems. These are variants of RNNs implementing the reservoir computing formalism, i.e. ANNs in which weights and biases are selected randomly. Andersson et al. (2019) used CNNs, which gained popularity in image recognition and allow for identifying patterns at different scales, while Ayed et al. (2019); Rahman et al. (2022) use ODE-nets, recently proposed as a new paradigm for neural state-space modelling. ODE-nets treat the layers of an ANN as intermediate time steps of a multistep ODE solver and are commonly trained using the adjoint method, a ubiquitous tool in data assimilation.
@@ 2.3 Model Based Reinforcement Learning and Optimal Control
MBRL methods can be classified based on how the model is built and how it is used. Concerning the approaches for model constructions, a large arsenal of function approximations has been deployed in the literature. ANNs (Hunt et al., 1992; Kurutach et al., 2018; Nagabandi et al., 2017), Gaussian Processes (GPr (Deisenroth and Rasmussen, 2011; Boedecker et al., 2014a)) and time-varying linear models (Boedecker et al., 2014b) are some examples. Concerning model usage, the simplest approach consists of using it solely to sample the state-action space, while more advanced approaches use the model in the training steps. One of the seminal MBRL approaches is the Dyna algorithm proposed by Sutton (1991). This uses experience (i.e. trajectories in the state-action space and associated rewards) collected from both environment and model. More sophisticated approaches use the model to “look ahead”, i.e. predicting the future evolution of the system under a given policy. A notable example is the Monte Carlo Tree Search in the AlphaGO (Silver et al., 2016, 2018) algorithm that reached superhuman performances in the game of Go. Moving from Markov Decision Processes to dynamical systems, the boundaries between MBRL and Model Predictive Control (MPC) become blurred, particularly when the latter uses black-box models (see Hedengren et al. 2014; Raw 2000 and Nagabandi et al. 2017 or Weber et al. 2017).
@@ 3 Definitions and General Formulation
  • The model-free policy ($\pi_{mf}$) training relies on classic model-free reinforcement learning. The optimal weights $\bm{w}_{a,mf}$ are defined according to a reward function, which measures how well the system is kept close to the desired trajectory. This evaluation is made online, updating the policy after each episode. The optimal weights maximize this function, which is written as the summation of instantaneous rewards $r_{k}=r(\check{\bm{e}}_{k},\check{\bm{a}}_{k})\in\mathbb{R}$, discounted by a factor $\gamma$:
    $\mathcal{R}_{c}(\bm{a})=\sum^{n_{t}-1}_{k=0}\gamma^{k}r_{k}(\check{\bm{e}}_{k},\check{\bm{a}}_{k})\,.$ (1)
@@
[…]
    with $\mathcal{L}_{p}:\mathbb{R}^{n_{s}\times n_{s}}\rightarrow\mathbb{R}$ the Lagrangian function for the identification problem and the expectation $\mathbb{E}_{\sim S(\bm{z},\bm{w}_{p})}\{\}$ over possible virtual trajectories accounting for the stochastic exogenous inputs. More specifically, starting from one sample signal $\bm{z}_{k}$, with $k=[0,n_{t}-1]$, we build a set of $N_{z}$ realizations of exogenous inputs as described in section 4.2 and we denote as $S(\bm{z},\bm{w}_{p})$ the set of trajectories generated for a given sample of the exogenous input $\bm{z}$ and a given set of weights $\bm{w}_{p}$. Two reasons motivate the choice for the integral formulation in (2) even if only a discrete set of samples $\hat{\bm{s}}(t_{k})$ is available. First, the integral is particularly handy in deriving the adjoint problem driving the computation of the gradient $d_{p}\mathcal{J}$, as presented in section 4.2. Second, the choice keeps the approach independent from the specific numerical solver to advance the virtual states $\bm{s}(t)$. We return to both points in section 4.3. We here stress that the closure model $g(\bm{s};\bm{w}_{g})$ makes the digital twin predictive only for a given sample of the exogenous disturbance signal and within $t\in[0,T_{o}]$, i.e. no forecasting is considered.
    The model-based control seeks to find the optimal policy to keep the digital twin along the desired trajectory. We introduce the virtual error $\bm{e}(t_{k})=e(\bm{s}_{k},\tilde{\bm{s}}_{k})$ to measure the distance between the current virtual state and the target state and $\bm{w}_{a,mb}=\pi_{mb}(\bm{e};\bm{w}_{a,mb})$ the virtual actions taken on the digital twin. The performances of the controller are measured by the cost function $\mathcal{J}_{a}(\bm{w}_{a,mb})$. For convenience, we define this cost function similarly to the one in (2), i.e.:
[…]
@@ 4.1 The Model-Free loop: (1)-(2)-(3)-(4)
We consider a classic actor-critic reinforcement learning formalism (Sutton and Barto 2018, Bhatnagar et al. 2009). Although more recent variants exist, we here focus on the classic Deep Deterministic Policy Gradient (DDPG) algorithm proposed by Lillicrap et al. (2019) with the prioritized experience replay strategy by Schaul et al. (2015). This algorithm was extensively benchmarked by Pino et al. (2023) in flow control problems and offers a good compromise between performance and simplicity.
The DDPG trains two ANNs while interacting with the system. The first, referred to as actor, aims to approximate the policy function $\pi_{mf}:\check{\bm{e}}\rightarrow\check{\bm{a}}$, which provides the action to be taken at a certain state. This policy is deterministic, but we add noise generated by a random process to explore action space.
The second, referred to as critic, aims to approximate the state action value function $q:(\check{\bm{e}},\check{\bm{a}})\rightarrow Q$; this provides the expected reward for taking action $\check{\bm{a}}$ while the system is in the error state $\check{\bm{e}}$ and then following a certain policy. We use $Q$ to denote the (unknown) true Q value and $q(\check{\bm{e}},\check{\bm{a}};\bm{w}_{q})$ to denote the approximation from the $Q-$network. Referring to the previous section and Figure 1, the parameters $\bm{w}_{a}$ denote the weights and biases of the actor-network while $\bm{w}_{q}$ denotes the weights and biases for the critic network.
@@
Finally, the sampling from the buffer is carried out in the form of batches to evaluate the gradient $d\mathcal{J}_{q}/d\bm{w}_{q}$. The sampling prioritizes transitions with the largest TD-error, $\delta_{k}$, (Schaul et al., 2016) building a non-zero probability distribution for each sample proportional to it $p_{i}=|\delta_{i}|+\mathrm{tol}$, yielding to the sampling probability of the $i-$th transition, $P(i)$ equal to:
@@ 4.2 The Model-Based Loop (1)- (2)- (3)-(5)-(6)
$\tilde{\bm{d}}[j](t;\bm{\alpha}[j])=\sum^{n_{t}-1}_{k=0}\alpha[j]_{k}\kappa(t_{k},t)$ (14)
@@ 5 Initialize Step 6 counters and thresholds $C_{\pi}$, $P_{\pi}$, $C_{w}$, $P_{w}$, $T_{J}$ and $T_{w}$
16             Sample $\textsc{min}(n_{b},n_{t})$ transitions from $\mathcal{D}_{\mathrm{L}}$ and compute $\mathcal{J}_{q}$ and $d\mathcal{J}_{q}/d\bm{w}_{q}$ from (7)
@@ 50       Policy switching and weight cloning based on the decision tree (Figure 3)
  • Step 2: Critic Update and model-free policy update (lines 14-24). An optimization with $n_{Q}$ iterations is carried out on the critic network, followed by an optimization with $n_{A}$ iterations for the policy. Both use batches of $n_{b}$ transitions sampled from $\mathcal{D}_{\mathrm{L}}$. Concerning the training of the critic network (lines 14-19), this loss function and its gradient are computed from (7). This step is essentially a training of the critic network in a supervised learning formalism. Similarly to what was proposed by Lillicrap et al. (2019) for the original DDPG, the update of the critic is modulated by an under-relaxation $\tau$ using a target network (line 17). At every $n_{Q}$ iteration, the optimized critic re-evaluates the TD errors of the sampled transitions in $\mathcal{D}_{\mathrm{L}}$ (line 18). Limiting the TD updates for the current batch of trajectories allows for a pseudo-random update of the most critical transitions, according to the critic. In addition, this approach is also computationally cheaper than re-evaluating all the stored transitions in the buffer, usually in the order of magnitude of $\mathcal{O}(10^{6})$. Successively, the critic is then used to train the policy in lines 20-24. In this second stage, the gradient with respect to the cost function is computed using eq (4). Note that we use the notation $\Delta\bm{w}$ to denote the optimization update for the weights $\bm{w}$, such that the updating reads $\bm{w}\leftarrow\bm{w}+\Delta\bm{w}$. Therefore, the update in a simple gradient descent over a cost function $J(\bm{w})$, for example, reads $\Delta\bm{w}=-\eta dJ(\bm{w})/d\bm{w})$. This makes the notation independent of the choice of the optimizer. We note that differently from the original DDPG algorithm in Lillicrap et al. (2019), the updates of the critic and the policy networks are carried out via two separate loops rather than in a single one. In our experiments, we found this approach to lead to more stable model-free policy updates.
@@
  • Step 4: Optimize digital twin on the real actions (lines 27-36). The goal of this step is to find the closure parameters $\bm{w}_{p}$ such that the digital twin’s prediction matches with the $n_{S}$ trajectories of the real system, stored in $\mathcal{D}_{\mathrm{S}}$. Consequently, the loss function $\mathcal{J}_{p}$ in (2) is minimized using a gradient-based optimization loop. This consists of $n_{G}$ iterations and computes the updates $\Delta\bm{w}_{p}$ from the gradient evaluated along the $n_{S}$ trajectories available in $\mathcal{D}_{\mathrm{S}}$. Again, fixing the numerical time stepping to the one of the real interaction avoids the need for action interpolation, but might become problematic for stiff problems. For each trajectory (loop in lines 27-33), a total of $N_{z}$ possible realizations of the exogenous input are considered (sampled from the processes $\mathcal{N}^{(l)}_{z}$ from the previous step), and the gradient is computed using the adjoint method (see eq. 10). The associated updates are then averaged (line 32), and the optimization step is carried out (line 34). Each optimization step requires a total of $n_{S}\times N_{Z}$ forward and backward virtual episodes.
@@
  • Step 6: Policy switching and parametrization cloning. (lines 44-47). This step decides which of the policies $\pi_{mb}$ or $\pi_{mf}$ becomes “live” and which one becomes “idle”. The first is deployed in the next interaction with the real system, while the second continues to be trained with different mechanisms. The decision in this step is based on the policy performances in the virtual environment, evaluated and averaged over $N$ episodes with randomized initial conditions and disturbances. Defining as $\bar{\mathcal{J}}_{a,\mathrm{live}}$ and $\bar{\mathcal{J}}_{a,\mathrm{idle}}$ the averaged cost functions for the “live” and “idle” policies, we introduce a tolerance factor $T_{J}\in\mathbb{R}$, a threshold $P_{\pi}\in\mathbb{N}$, and a counter $C_{\pi}\in\mathbb{N}$ to avoid overly impulsive decisions. If the live policy performs worse than the idle one more than the tolerance factor ($\bar{\mathcal{J}}_{a,\mathrm{live}}/\bar{\mathcal{J}}_{a,\mathrm{idle}}<T_{J}$), the counter $C_{\pi}$ gets incremented. If $C_{\pi}$ exceeds the threshold $P_{\pi}$, the idle policy takes over and becomes the live one. On the other hand, if the opposite is true, then we check that the idle policy did not happen to be stuck in a local minima, analyzing its weight variation:
@@
    where $i$ indicates the current number of the episode ($i\in[1,n_{e}]$). To this end, we introduce a tolerance $T_{w}$, a counter $C_{w}$ and a threshold $P_{w}$ and repeat the previous decision process: if $d\bm{w}_{a}$ falls below $T_{w}$, $C_{w}$ increases. If $C_{w}$ goes beyond $P_{w}$, it is reasonable to assume that the policy optimization step is trapped in a local minimum, and we replace the “idle” parametrization by cloning the “live” one. The decision tree of this step is recalled in Figure 3.
@@
In Steps 4 and 5, the $n_{S}$ hyperparameters define how much the digital twin should consider past experiences. In this work, we set $n_{S}$=1, thus focusing on the last trajectory solely, while having a variable $N_{z}$ to promote robustness against randomness on the exogenous inputs and take into account possible sampling errors which may occur in real-world scenarios.

### 25 Feb 2024 v3

@@ Abstract
The concept of digital twins promises to revolutionize engineering by offering new avenues for optimization, control, and predictive maintenance. We propose a novel framework for simultaneously training the digital twin of an engineering system and an associated control agent. The training of the twin combines methods from adjoint-based data assimilation and system identification, while the training of the control agent combines model-based optimal control and model-free reinforcement learning. The training of the control agent is achieved by letting it evolve independently along two paths: one driven by a model-based optimal control and another driven by reinforcement learning. The virtual environment offered by the digital twin is used as a playground for confrontation and indirect interaction. This interaction occurs as an “expert demonstrator”, where the best policy is selected for the interaction with the real environment and “cloned” to the other if the independent training stagnates. We refer to this framework as Reinforcement Twinning (RT).
@@ 1 Introduction
Nonlinear system identification is mostly concerned with the model identification problem. The system is subject to external inputs and observed through noisy measurements but is deterministic. The model can be fit into structures that are derived either from first principles or from general templates such as NRAX or Volterra models (Schoukens and Ljung, 2019), and the main goal is to predict how the system responds to actuation. The traditional context is control engineering, i.e. driving the system towards a desired trajectory. Besides the stronger focus on input-output relation (and thus the focus on actuated systems), the main difference with respect to traditional data assimilation is in the deterministic nature of the system to be inferred: errors in the initial conditions are not amplified as dramatically as, for example, in weather forecasting. Therefore, the resulting uncertainties remain within the bounds of model uncertainties and the state estimation is less critical than in data assimilation. Moreover, in many actuated systems, the forced response dominates over the natural one and the initial conditions are quickly forgotten. Tools for system identification are reviewed by Nelles (2001) and Suykens et al. (1996).
@@
While machine learning enters the literature of data assimilation and system identification, cross-fertilization proceeds bilaterally. The need for identifying dynamic models from data is also growing in the literature on reinforcement learning, which is significantly drawing ideas from system identification and optimal control theory. Reinforcement learning is a subset of machine learning concerned with the training of an agent via trial and error to achieve a goal while acting on an environment (Sutton and Barto, 2018; Bertsekas, 2019). The framework differs from classic control theory because of its roots in sequential decision making: the environment to be controlled and the agent constitute a Markov Decision Process (MDP, Puterman (1994)) rather than a dynamical system with the usual state-space representation. Model-free approaches solely rely on input-output information and function approximations of the agent’s performance measures, while model-based approaches combine these with a model of the environment (Moerland et al., 2022a; Luo et al., 2022a). The model can be used as a forecasting tool, thus playing the same role as in model predictive control, or as a playground to accelerate the learning (Schwenzer et al., 2021). Model-free approaches have gained burgeoning popularity thanks to their success in video games (Szita, 2012; Mnih et al., 2013a) or natural language processing (Uc-Cetina et al., 2022), for which the definition of a system model is cumbersome. However, these algorithms have shown significant and somewhat surprising limitations for engineering problems governed by simple PDEs (Werner and Peitz, 2023; Pino et al., 2023).
This article proposes a framework combining ideas from data assimilation, system identification and reinforcement learning with the goal of: (1) training a digital twin on real-time data and (2) solving a control problem. We refer to this framework as Reinforcement Twinning (RT). The rest of the article is structured as follows. Section 2 provides a concise literature review of related works across the intersected disciplines and identifies the main novelties of the proposed framework. Section 3 presents the general framework and the key definitions, while Section 4 goes into the mathematical details of its implementation. Section 5 introduces three test cases selected for their relevance in the engineering literature and stemming from largely different fields. These are (1) the control of a wind turbine subject to time-varying wind speed, (2) the trajectory control of flapping-wing micro air vehicles (FWMAVs) subject to wind gusts, and (3) the mitigation of thermal loads in the management of cryogenic storage tanks. Although all test cases are analyzed using synthetic data that might look somewhat contrived, the main focus is here to show that the model-based and the model-free training can benefit from each other. Section 6 presents the main results; section 7 closes with conclusions and perspectives.
@@ 2.2 Machine learning for System Identification and Control
Important contributions within the first class of approaches are the works by Ljung et al. (2020), Gonzalez and Yu (2018) and Bucci et al. (2018) who use LSTM networks (Hochreiter and Schmidhuber, 1997). These are variants of RNNs, which are very popular in natural language processing and use three gates (forget/input/output) to preserve information over longer time steps, thus allowing for better handling of long-term dependencies. Canaday et al. (2020) used ESNs, which gained popularity for forecasting in stochastic systems. These are variants of RNNs implementing the reservoir computing formalism, i.e. ANNs in which weights and biases are selected randomly. Andersson et al. (2019) used CNNs, which gained popularity in image recognition and allow for identifying patterns at different scales, while Ayed et al. (2019) and Rahman et al. (2022) use ODE-nets, recently proposed as a new paradigm for neural state-space modelling. ODE-nets treat the layers of an ANN as intermediate time steps of a multistep ODE solver and are commonly trained using the adjoint method, a ubiquitous tool in data assimilation. An overview of the challenges of learning models of dynamical systems from data has been recently presented by Bucci et al. (2023), who also propose a “gradual” approach which seeks to build models of gradually increasing complexity.
@@ 2.3 Model Based Reinforcement Learning and Optimal Control
MBRL methods can be classified based on how the model is built and how it is used. Concerning the approaches for model constructions, a large arsenal of function approximations has been deployed in the literature. ANNs (Hunt et al., 1992; Kurutach et al., 2018; Nagabandi et al., 2017), Gaussian Processes (GPr,Deisenroth and Rasmussen (2011); Boedecker et al. (2014a)) and time-varying linear models (Boedecker et al., 2014b) are some examples. Concerning model usage, the simplest approach consists of using it solely to sample the state-action space, while more advanced approaches use the model in the training steps. One of the seminal MBRL approaches is the Dyna algorithm proposed by Sutton (1991). This uses experience (i.e. trajectories in the state-action space and associated rewards) collected from both environment and model. More sophisticated approaches use the model to “look ahead”, i.e. predicting the future evolution of the system under a given policy. A notable example is the Monte Carlo Tree Search in the AlphaGO (Silver et al., 2016, 2018) algorithm that reached superhuman performances in the game of Go. Moving from Markov Decision Processes to dynamical systems, the boundaries between MBRL and Model Predictive Control (MPC) become blurred, particularly when the latter uses black-box models (see Hedengren et al. 2014; Raw 2000 and Nagabandi et al. 2017 or Weber et al. 2017).
@@ 3 Definitions and General Formulation
  • The model-free policy ($\pi_{mf}$) training relies on classic model-free reinforcement learning. The optimal weights $\bm{w}_{a,mf}$ are defined according to a reward function, which measures how well the system is kept close to the desired trajectory. This evaluation is made online, updating the policy after each episode. The optimal weights maximize this function, which is written as the summation of instantaneous rewards $r_{k}=r(\check{\bm{e}}_{k},\check{\bm{a}}_{k})\in\mathbb{R}$, possibly discounted by a factor $\gamma$:
    $\mathcal{R}_{c}(\bm{a},\bm{e})=\sum^{n_{t}-1}_{k=0}\gamma^{k}r_{k}(\check{\bm{e}}_{k},\check{\bm{a}}_{k})\,.$ (1)
@@
    $\mathcal{J}_{p}(\bm{w}_{p})=\mathbb{E}_{\sim S(\bm{z},\bm{w}_{p})}\biggl{\{}\int^{T_{o}}_{0}\mathcal{L}_{p}\big{(}\check{\bm{s}}(t_{k}),\bm{s}(t;\bm{w}_{p})\big{)}\;\mathrm{dt}\,\biggr{\}},$ (2)
    with $\mathcal{L}_{p}:\mathbb{R}^{n_{s}\times n_{s}}\rightarrow\mathbb{R}$ the Lagrangian function for the identification problem and the expectation $\mathbb{E}_{\sim S(\bm{z},\bm{w}_{p})}\{\}$ over possible virtual trajectories accounting for the stochastic exogenous inputs. More specifically, starting from one sample signal $\bm{z}_{k}$, with $k=0,n_{t}-1$, we build a set of $N_{z}$ realizations of exogenous inputs as described in section 4.2 and we denote as $S(\bm{z},\bm{w}_{p})$ the set of trajectories generated for a given sample of the exogenous input $\bm{z}$ and a given set of weights $\bm{w}_{p}$. Two reasons motivate the choice for the integral formulation in (2) even if only a discrete set of samples $\hat{\bm{s}}(t_{k})$ is available. First, the integral is particularly handy in deriving the adjoint problem driving the computation of the gradient $d\mathcal{J}/d\bm{p}$, as presented in section 4.2. Second, the choice keeps the approach independent from the specific numerical solver to advance the virtual states $\bm{s}(t)$. We return to both points in section 4.3. We here stress that the closure model $g(\bm{s};\bm{w}_{g})$ makes the digital twin predictive only for a given sample of the exogenous disturbance signal and within $t\in[0,T_{o}]$, i.e. no forecasting is considered.
    The model-based control seeks to find the optimal policy to keep the digital twin along the desired trajectory. We introduce the virtual error $\bm{e}(t_{k})=e(\bm{s}_{k},\tilde{\bm{s}}_{k})$ to measure the distance between the current virtual state and the target state and $\bm{a}_{mb}=\pi_{mb}(\bm{e};\bm{w}_{a,mb})$ the virtual actions taken on the digital twin. The performances of the controller are measured by the cost function $\mathcal{J}_{a}(\bm{w}_{a,mb})$. For convenience, we define this cost function similarly to the one in (2), i.e.:
    $\mathcal{J}_{a}(\bm{w}_{a,mb})=\mathbb{E}_{\sim S(\bm{z},\bm{w}_{a,mb}}\biggl{\{}\int^{T_{o}}_{0}\mathcal{L}_{a}\big{(}\tilde{\bm{s}}(t),\bm{s}(t;\bm{w}_{a,mb})\big{)}\;\mathrm{dt}\,\biggr{\}}\,,$ (3)
@@ 4.1 The Model-Free loop: (1)-(2)-(3)-(4)
We consider a classic actor-critic reinforcement learning formalism (Sutton and Barto 2018, Bhatnagar et al. 2009). Although more recent variants exist, we here focus on the classic Deep Deterministic Policy Gradient (DDPG) algorithm proposed by Lillicrap et al. (2019) with the prioritized experience replay strategy by Schaul et al. (2015). This algorithm offers a good compromise between performance and simplicity, and it has been proven successful in several flow control tasks (Bucci et al., 2019; Pino et al., 2023).
The DDPG trains two functions while interacting with the system. The first, referred to as actor, aims to approximate the policy function $\pi_{mf}:\check{\bm{e}}\rightarrow\check{\bm{a}}$, which provides the action to be taken at a certain state. This policy is deterministic, but we add noise generated by a random process to explore action space.
The second, referred to as critic, aims to approximate the state action value function $q:(\check{\bm{e}},\check{\bm{a}})\rightarrow Q$; this provides the expected reward for taking action $\check{\bm{a}}$ while the system is in the error state $\check{\bm{e}}$ and then following a certain policy. We use $Q$ to denote the (unknown) true Q value and $q(\check{\bm{e}},\check{\bm{a}};\bm{w}_{q})$ to denote the approximation from the $Q-$network. Referring to the previous section and Figure 1, the parameters $\bm{w}_{a,mf}$ denote the weights and biases of the actor-network while $\bm{w}_{q}$ denotes the weights and biases for the critic network.
@@
Finally, the sampling from the buffer is carried out in the form of batches to evaluate the gradient $d\mathcal{J}_{q}/d\bm{w}_{q}$. The sampling prioritizes transitions with the largest TD-error, $\delta_{k}$, (Schaul et al., 2016) building a non-zero probability distribution for each sample proportional to it $p_{i}=|\delta_{i}|+\mathrm{tol}$, indicating with tol a small offset. This yields to the sampling probability of the $i-$th transition, $P(i)$ to be expressed as:
@@ 4.2 The Model-Based Loop (1)- (2)- (3)-(5)-(6)
$\tilde{\bm{d}}[j](t;\bm{\alpha}[j])=\sum^{n_{t}-1}_{k=0}\kappa(t_{k},t)\alpha[j]_{k}$ (14)
@@ 5 Initialize Step 6 counters and thresholds $C_{\pi}$, $P_{\pi}$, $C_{w}$, $P_{w}$, $T_{J}$ and $T_{w}$
16             Sample $\textsc{min}(n_{b},n_{t})$ transitions from $\mathcal{D}_{\mathrm{L}}$ and compute $\mathcal{J}_{q}$ and $d\mathcal{J}_{q}/dq$ from (7)
@@ 50       Policy switching and weight cloning based on the decision tree (Figure 3)
  • Step 2: Critic Update and model-free policy update (lines 14-24). An optimization with $n_{Q}$ iterations is carried out on the critic network, followed by an optimization with $n_{A}$ iterations for the policy. Both use batches of $n_{b}$ transitions sampled from $\mathcal{D}_{\mathrm{L}}$. Concerning the training of the critic network (lines 14-19), this loss function and its gradient are computed from (7). This step is essentially a training of the critic network in a supervised learning formalism. Similarly to what was proposed by Lillicrap et al. (2019) for the original DDPG, the update of the critic is modulated by an under-relaxation $\tau$ using a target network (line 17). At every $n_{Q}$ iteration, the optimized critic re-evaluates the TD errors of the sampled transitions in $\mathcal{D}_{\mathrm{L}}$ (line 18). Limiting the TD updates for the current batch of trajectories allows for a pseudo-random update of the most critical transitions, according to the critic. In addition, this approach is also computationally cheaper than re-evaluating all the stored transitions in the buffer, usually in the order $\mathcal{O}(10^{6})$. Successively, the critic is then used to train the policy in lines 20-24. In this second stage, the gradient with respect to the cost function is computed using eq (4). Note that we use the notation $\Delta\bm{w}$ to denote the optimization update for the weights $\bm{w}$, such that the updating reads $\bm{w}\leftarrow\bm{w}+\Delta\bm{w}$. Therefore, the update in a simple gradient descent over a cost function $J(\bm{w})$, for example, reads $\Delta\bm{w}=-\eta\;dJ(\bm{w})/d\bm{w}$. This makes the notation independent of the choice of the optimizer. Differently from the original DDPG algorithm in Lillicrap et al. (2019), the updates of the critic and the policy networks are carried out via two separate loops rather than in a single one. In our numerical experiments, we found this approach to lead to more stable model-free policy updates.
@@
  • Step 4: Optimize digital twin on the real actions (lines 27-36). The goal of this step is to find the closure parameters $\bm{w}_{p}$ such that the digital twin’s prediction matches with the $n_{S}$ trajectories of the real system, stored in $\mathcal{D}_{\mathrm{S}}$. Consequently, the loss function $\mathcal{J}_{p}$ in (2) is minimized using a gradient-based optimization loop. This consists of $n_{G}$ iterations and computes the updates $\Delta\bm{w}_{p}$ from the gradient evaluated along the $n_{S}$ trajectories available in $\mathcal{D}_{\mathrm{S}}$. Again, fixing the numerical time stepping to the one of the real interaction avoids the need for action interpolation, but might become problematic for stiff problems. For each trajectory (loop in lines 27-33), a total of $N_{z}$ possible realizations of the exogenous input are considered (sampled from the $\mathcal{N}^{(l)}_{z}$ processes built in the previous step), and the gradient is computed using the adjoint method (see eq. 10). The associated updates are then averaged (line 32), and the optimization step is carried out (line 34). Each optimization step requires a total of $n_{S}\times N_{Z}$ forward and backward virtual episodes.
@@
  • Step 6: Policy switching and parametrization cloning. (lines 44-47). This step decides which of the policies $\pi_{mb}$ or $\pi_{mf}$ becomes “live” and which one becomes “idle”. The first is deployed in the next interaction with the real system, while the second continues to be trained with different mechanisms. The decision in this step is based on the policy performances in the virtual environment, evaluated and averaged over $N$ episodes with randomized initial conditions and disturbances. The flow chart for this step is illustrated in Figure 3. Defining as $\bar{\mathcal{J}}_{a,\mathrm{live}}$ and $\bar{\mathcal{J}}_{a,\mathrm{idle}}$ the averaged cost functions for the “live” and “idle” policies, we introduce a tolerance factor $T_{J}\in\mathbb{R}$, a threshold $P_{\pi}\in\mathbb{N}$, and a counter $C_{\pi}\in\mathbb{N}$ to avoid overly impulsive decisions. If the live policy performs worse than the idle one more than the tolerance factor ($\bar{\mathcal{J}}_{a,\mathrm{live}}/\bar{\mathcal{J}}_{a,\mathrm{idle}}<T_{J}$), the counter $C_{\pi}$ gets incremented. If $C_{\pi}$ exceeds the threshold $P_{\pi}$, the “idle” policy takes over and becomes “live”. On the other hand, if the opposite is true, we check that the idle policy is not stuck in a local minima, by analyzing its weight variation:
@@
    where $i$ indicates the current number of the episode ($i\in[1,n_{e}]$). To this end, we introduce a tolerance $T_{w}$, a counter $C_{w}$ and a threshold $P_{w}$ and repeat the previous decision process: if $d\bm{w}_{a}$ falls below $T_{w}$, $C_{w}$ increases. If $C_{w}$ goes beyond $P_{w}$, it is reasonable to assume that the policy optimization step is trapped in a local minimum, and we replace the “idle” parametrization by cloning the “live” one.
@@
Tthe $n_{S}$ hyperparameter in Steps 4 and 5 defines how much the digital twin should consider past experiences. In this work, we set $n_{S}$=1, thus focusing on the last trajectory solely, while having a variable $N_{z}$ to promote robustness against randomness on the exogenous inputs and take into account possible sampling errors which may occur in real-world scenarios.